A statistics course has seven male and three female students. The professor wants to select two students at random to help her conduct a research project. What is the probability that the two students chosen are female?


1
Expert's answer
2022-03-16T14:52:47-0400

P(both female) = "{\frac {3 \choose 2} {10 \choose 2}}={\frac 3 {120}}={\frac 1 {40}}"